OpenStudy (anonymous):

Put the following equation in slope-intercept form: y - 5 = (x - 3)

hartnn (hartnn):

add 5 to both sides, what u get ?

hartnn (hartnn):

*slope-intercept form is y=mx+c

OpenStudy (anonymous):

sorry its actually y-5=2/7(x-3)

hartnn (hartnn):

no problem, add 5 to both sides, what u get ?

OpenStudy (anonymous):

y=5+2/7(x-3)?

hartnn (hartnn):

now, yo need to simplify right side, (2/7) (x-3) =.... ? use a(b-c) = ab-ac ...that is multiply them out...

OpenStudy (anonymous):

2/7x-6/7

hartnn (hartnn):

yes, now y = 5+ 2/7x-6/7 you just need to smplify 5-6/7 can u

OpenStudy (anonymous):

....noo??

hartnn (hartnn):